# Learning-Augmented Skip Lists

### v1

@@ 4.3 AOL Dataset
Fig. 6: AOL datasets distribution characterization. This figure illustrates how the $\alpha$ value of 0.75 is obtained for the AOL dataset. The AOL dataset shows a much smaller $\alpha$ value compared to the CAIDA dataset so AOL almost resembles a uniform distribution despite very few high-frequency nodes. This also explains why the performance of the augmented skip list is close to the vanilla implementation.
@@
Fig. 7: Insertion and query time on AOL of vanilla and augmented skip lists
5 Conclusion
In this paper, we studied learning-augmented algorithms for data structure design. In particular, we incorporated the learning-augmented paradigm into improving upon the traditional skip-list data structures. We showed that given reasonably accurate advice on the ubiquitous Zipfian distribution, our data structure achieves significantly better performance than traditional skip lists. Moreover, our data structure achieves both consistency and robustness, so that our algorithm does not overly suffer poor performance in the case that the advice is inaccurate. We showed experimentally on both synthetic and real-world datasets that our learning-augmented skip list performs at least as well as a traditional random skip list, irrespective of the oracle’s accuracy and the nature of the input distribution. Looking forward, there is potential for further exploration into skip list implementation that responds dynamically to dynamic datasets. We hope our work inspires future studies to incorporate machine learning heuristics into provably better data structures.

### v2

@@ D.1.2 AOL Dataset
Figure 9: AOL datasets distribution characterization. This figure illustrates how the $\alpha$ value of 0.75 is obtained for the AOL dataset. The AOL dataset shows a much smaller $\alpha$ value compared to the CAIDA dataset so AOL almost resembles a uniform distribution despite very few high-frequency nodes. This also explains why the performance of the augmented skip list is close to the classic implementation.
@@
Figure 10: Insertion and query time on AOL of classic and augmented skip lists
D.2 KD Trees
For KD Trees, first describe our methodology for evaluating our data structure on synthetic data. We then describe our empirical evaluations on real-world datasets.
The computer used for KD tree benchmarking is a desktop machine with an Intel Core i9-14900KF@ 3200MHz, with 64GB RAM, and 2TB of Solid State Drive. The tests were conducted in Windows 10 Enterprise, version 10.0.19045 Build 19045.
D.2.1 Synthetic Datasets with Perfect Knowledge
First, consider a dataset with a Zipfian distribution. In order to construct this dataset, we first select $n$ unique data points in $[\Delta]^{d}$ uniformly. We then generate Zipfian frequencies, $f_{i}\approx\frac{1}{(i+b)^{a}}$, and randomly pair the frequencies to the data points to serve as both the construction and query frequencies. We construct either a traditional KD tree, or our learned KD tree on this dataset. Then, we evaluate the performance of querying by sampling the known datapoints with probabilities given by their Zipfian probabilities.
In Table 3, we use a Zipfian distribution with parameters $a=1$ and $b=2.7$. This data demonstrates that our method outperforms a traditional KD tree, given that the Zipfian distribution has the given parameters. In Figure 11, we vary the Zipfian parameters of our dataset. As expected, our learned KD tree performance increases with how skew the distribution is. Moreover, we find that our method outperforms the traditional KD tree on all tested Zipfian distributions.
type dim const_time avg query(s) avg query depth
traditional 1 1.267E-01 2.645E-05 1.330E+01
learned 1 2.387E-01 2.181E-05 1.086E+01
traditional 2 1.266E-01 2.709E-05 1.341E+01
learned 2 3.232E-01 2.221E-05 1.086E+01
traditional 3 1.253E-01 2.700E-05 1.334E+01
learned 3 3.981E-01 2.264E-05 1.097E+01
traditional 4 1.185E-01 2.724E-05 1.336E+01
learned 4 4.549E-01 2.256E-05 1.089E+01
traditional 5 1.300E-01 2.743E-05 1.336E+01
learned 5 5.286E-01 2.296E-05 1.096E+01
traditional 10 1.277E-01 2.896E-05 1.340E+01
learned 10 8.642E-01 2.403E-05 1.091E+01
traditional 20 1.295E-01 3.121E-05 1.344E+01
learned 20 1.543E+00 2.564E-05 1.085E+01
traditional 40 1.361E-01 3.558E-05 1.340E+01
learned 40 2.911E+00 2.902E-05 1.078E+01
Table 3: We construct and query KD trees with our method and with a traditional KD tree on synthetic datasets of various dimensionality. We construct our tree on 10k points, and query 1M times. We find that, independent of the data dimensionality, our method produces lower average query depths.
Figure 11: We generate datasets of $2^{12}$ points in 3-dimentional space, with frequencies given by a Zipfian distribution with parameters $a,b$. We then query the tree $2^{14}$ times, with point queries selected by the same Zipfian distribution. We repeat this process 32 times, and report the median of the average query depth across all runs. We find that our method is able to outperform the traditional KD tree method across all tested Zipfian distributions. Moreover, our performances increases as $a$ increases and $b$ decreases, making the Zipfian distribution most skew. Notably, when $a=0$, all points have uniform weights, at which point our method performs equivalently to the traditional KD tree.
D.2.2 Synthetic Datasets with Noisy Knowledge
In reality, it is rarely the case that we have perfect knowledge in constructing a model. In order to evaluate the performance of our method on noisy data, we create a synthetic dataset with noisy training information.
As before, we generate world points and assign them Zipfian weights. When constructing the tree, each weight us updated to be $Mp_{i}+A$, where $M$ and $A$ are drawn from uniform distributions. This new noisy distribution is normalized in order to form a valid probability distribution. Then, points are queried many times with their ground truth Zipfian probabilities. For the fixed Zipfian distribution with $a=5,b=2$, we plot the effects of different ranges of $M$ and $A$ to demonstrate the effect noise has in Fig. 4. This figure demonstrates that, even with moderate amounts of noise, our method still outperforms a traditional KD tree. Moreover, our method still remains on par with the traditional KD tree when significant noise is present, due to our robustness guarantees.
D.2.3 Real-World Datasets
In addition to evaluating our method on synthetic data, we also evaluate results on real-world data.
First, we consider $n$-grams in various languages. We test on a pre-processed subset of the Google N-Gram dataset [Goo22, Goo12].
In order to evaluate our method, we convert an $n$-gram to a vector in $\mathbb{Z}^{n}$ with each entry indexing the words in the $n$-gram. We construct the learning-augmented and traditional KD trees, and show the performance of lookup with queries weighted by their ground truth frequency in Table 4. In all cases, we find that the learning-augmented KD tree outperforms the traditional KD tree at average lookup depth.
Additionally, we test our method on a dataset of neuron activity, as provided by [ACL14], which has shown to be Zipfian. This dataset consists of vectors in $\{0,1\}^{30}$, indicating which of 30 cells fire at agiven time. As in their work, we bin in 20ms increments when constructing these vectors. We similarly ignore the time of observations, and build our learning-augmented KD tree with frequencies given by the rate of appearance of vectors. We find that, when querying with probabilities equivalent to the training distribution, a traditional KD tree has an average query depth of 23.7. Using our learning-augmented KD tree, however, we are able to achieve an average query depth of 14.9.
Dataset Traditional Avg Query Depth Learned Avg Query Depth
2grams_chinese_simplified 14.8388 12.1144
2grams_english-fiction 14.216 11.6232
2grams_english 14.7253 11.6647
2grams_french 14.5123 11.5999
2grams_german 13.9473 11.7066
2grams_hebrew 13.5322 12.0007
2grams_italian 14.5231 11.8027
2grams_russian 14.5769 11.8082
2grams_spanish 15.1763 11.6582
3grams_chinese_simplified 12.343 11.4211
3grams_english-fiction 13.5264 11.3183
3grams_english 15.002 11.3632
3grams_french 14.9235 11.3529
3grams_german 14.129 11.4285
3grams_hebrew 13.0887 9.6663
3grams_italian 13.3839 11.3616
3grams_russian 15.2475 11.1415
3grams_spanish 15.3855 11.3635
4grams_chinese_simplified 11.5823 9.8661
4grams_english-fiction 12.5141 9.7589
4grams_english 13.4464 9.7793
4grams_french 12.6383 9.835
4grams_german 12.3911 9.8765
4grams_hebrew 9.3198 7.4493
4grams_italian 11.3456 9.6834
4grams_russian 13.2406 9.6274
4grams_spanish 12.9712 9.8191
5grams_chinese_simplified 11.5373 9.8982
5grams_english-fiction 12.8027 9.8069
5grams_english 12.5607 9.5967
5grams_french 12.402 9.835
5grams_german 11.6912 9.8432
5grams_hebrew 7.097 6.2143
5grams_italian 11.3061 9.7771
5grams_russian 12.885 9.6805
5grams_spanish 12.111 9.8079
Table 4: We construct traditional and learning-augmented KD trees for $n$-grams for various languages, and of various lengths $n$. Our method outperforms a traditional KD tree in all cases.